OpenStudy (anonymous):

integral of xcos4x with explanation please?

OpenStudy (earthcitizen):

use substitution

myininaya (myininaya):

\[\int\limits_{}^{}x \sin(4x) dx=x \cdot \frac{-1}{4} \cos(4x)-\int\limits_{}^{}\frac{-1}{4} \cos(4x) dx\]

myininaya (myininaya):

integration by parts

OpenStudy (anonymous):

Is there any calculus problem that myin can't solve?

OpenStudy (earthcitizen):

are u sure it's not 4x^2

OpenStudy (earthcitizen):

myin, it's an indefinite integral

OpenStudy (anonymous):

\[1/4*\sin 4x *x +1/16*\cos 4x \]

myininaya (myininaya):

i know it is an indefinite integral integration by parts is the way to do this one the way i showed above

OpenStudy (earthcitizen):

*thunder*

OpenStudy (anonymous):

\[\int\limits_{}^{}x \sin 4xdx\]\[u = x, du = dx\]\[dv = sen(4x)dx\]\[v = (1/4)\cos (4x)\]\[\int\limits_{}^{}udv = uv - \int\limits_{}^{}vdu\]\[\int\limits_{}^{}x \sin 4xdx = x(1/4)\cos 4x - \int\limits_{}^{}(1/4)\cos(4x)dx\]= x(1/4)cos 4x - (1/4)(-(1/4)sen(4x)) =x(1/4)cos 4x + (1/16)sen(4x)